Identify the operations (addition, subtraction, division, multiplication) informing the following expressions and tell how the expressions have been formed.
 z + 1, z – 1, y + 17, y – 17

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

z+ 1 = 1 is added to z = Addition
z–1 = 1 is subtracted from z = Subtraction
y + 17 = 17 is added to y = Addition
y–17 = 17 is subtracted from y = Subtraction

​​Thanks